OpenStudy (anonymous):

The function y=Ce^kt passes through the two points: (1,4) and (4,19). Find C and k.

OpenStudy (anonymous):

HI :DDDD

OpenStudy (notamathgenius):

Hold On one second :)

OpenStudy (anonymous):

no problem:) I get what im supposed to do but I keep getting the problem wrong:/

OpenStudy (notamathgenius):

Hm, I'm not sure either >_< I'm sorry ;-;

OpenStudy (anonymous):

hmm its alright;) I'll just wait or ask a friend:)

OpenStudy (notamathgenius):

Alrighty, I hope you find the answer soon!!!! :) good luck

OpenStudy (anonymous):

Are you sure about that first point?

OpenStudy (anonymous):

oops its supposed to be (1,4)

OpenStudy (anonymous):

Okay, so the curve is given by \(y=Ce^{kt}\). You're given two points that are known to lie on the curve, \((1,4)\) and \((4,19)\). This means you have the following equations: \[\begin{cases}4=Ce^k\\19=Ce^{4k}\end{cases}\] In other words, when \(t=1\) you get \(y=4\). Similarly, when \(t=4\) you get \(y=19\). Solve for \(C\) and \(k\).

OpenStudy (anonymous):

Do I have to set the two equations equal to each other and find k first and then plug in the k into one of the equations to find C? I tried that and I kept getting the answer wrong:/ I think its probably how I did the math maybe..

OpenStudy (anonymous):

First thing I'd do would be to find an expression for one of the variables in terms of the other. So you could solve for \(C\) and get \(C=\dfrac{4}{e^k}=4e^{-k}\), as an example. Continuing in this manner, you would plug this into the other equation: \[19=Ce^{4k}~~\Rightarrow~~19=4e^{-k}\times e^{4k}=4e^{3k}\\ k=\cdots\] Then solve for \(C\).

OpenStudy (anonymous):

hmm, ill try that method, thanks.I'll see if I can get the correct answer this time:)